At 12 cm in the radial direction from a conductive sphere of radius 10 cm, the electric field is 1800 N/C and its orientation points towards the interior of the sphere. Find the surface density of the charge on the sphere.

1
Expert's answer
2022-03-01T17:32:48-0500

"E=\frac{kqr}{R^3},"

"q=\sigma S=4\pi \sigma R^2,"

"E=\frac{4\pi kr\sigma}R,"

"\sigma=\frac{ER}{4\pi kr}=0.24~\frac{\mu C}{m^2}."